\begin{thebibliography}{10}

\bibitem{AlevDumas:QuantumGK}
{\sc J.~Alev and F.~Dumas}, {\em Sur le corps des fractions de certaines
  alg\`ebres quantiques}, J. Algebra, 170 (1994), pp.~229--265.

\bibitem{AOV:CounterExamplesToGK}
{\sc J.~Alev, A.~Ooms, and M.~Van~den Bergh}, {\em A class of counterexamples
  to the {G}elfand-{K}irillov conjecture}, Trans. Amer. Math. Soc., 348 (1996),
  pp.~1709--1716.

\bibitem{AOV:GKUptoDimension8}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em The
  {G}elfand-{K}irillov conjecture for {L}ie algebras of dimension at most
  eight}, J. Algebra, 227 (2000), pp.~549--581.

\bibitem{Bavula:GWAI}
{\sc V.~V. Bavula}, {\em Generalized {W}eyl algebras and their
  representations}, Algebra i Analiz, 4 (1992), pp.~75--97.

\bibitem{Bavula:GlobalDim}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Tensor homological
  minimal algebras, global dimension of the tensor product of algebras and of
  generalized {W}eyl algebras}, Bull. Sci. Math., 120 (1996), pp.~293--335.

\bibitem{Caldero:quantumGelfandKirillov}
{\sc P.~Caldero}, {\em On the {G}elfand-{K}irillov conjecture for quantum
  algebras}, Proc. Amer. Math. Soc., 128 (2000), pp.~943--951.

\bibitem{CKPR11}
{\sc J.-L. Colliot-Th\'{e}l\`ene, B.~Kunyavski\u{\i}, V.~Popov, and
  Z.~Reichstein}, {\em Is the function field of a reductive {L}ie algebra
  purely transcendental over the field of invariants for the adjoint action?},
  Compos. Math., 147 (2011), pp.~428--466.

\bibitem{Millet:QuantumGK}
{\sc F.~Fauquant-Millet}, {\em Quantification de la localisation de {D}ixmier
  de {$U({\rm sl}_{n+1}({\bf C}))$}}, J. Algebra, 218 (1999), pp.~93--116.

\bibitem{FutornyHartwig:QuantumGK}
{\sc V.~Futorny and J.~Hartwig}, {\em Solution of a {$q$}-difference {N}oether
  problem and the quantum {G}elfand-{K}irillov conjecture for
  {$\mathfrak{gl}_N$}}, Math. Z., 276 (2014), pp.~1--37.

\bibitem{GelfandKirillov:Conjecture-I}
{\sc I.~M. Gelfand and A.~A. Kirillov}, {\em Fields associated with enveloping
  algebras of {L}ie algebras}, Soviet Math. Doklady 7,  (1966), pp.~407--409.

\bibitem{GelfandKirillov:Conjecture-II}
\leavevmode\vrule height 2pt depth -1.6pt width 23pt, {\em Sur les corps
  li\'{e}s aux alg\`ebres enveloppantes des alg\`ebres de {L}ie}, Inst. Hautes
  \'{E}tudes Sci. Publ. Math.,  (1966), pp.~5--19.

\bibitem{Hodges:NCDeformations}
{\sc T.~J. Hodges}, {\em Noncommutative deformations of type-{$A$} {K}leinian
  singularities}, J. Algebra, 161 (1993), pp.~271--290.

\bibitem{IoharaMalikov:QuantumGK}
{\sc K.~Iohara and F.~Malikov}, {\em Rings of skew polynomials and
  {G}el'fand-{K}irillov conjecture for quantum groups}, Comm. Math. Phys., 164
  (1994), pp.~217--237.

\bibitem{Joseph:GKConjecture}
{\sc A.~Joseph}, {\em Proof of the {G}elfand-{K}irillov conjecture for solvable
  {L}ie algebras}, Proc. Amer. Math. Soc., 45 (1974), pp.~1--10.

\bibitem{kaygun:GWAs}
{\sc A.~Kaygun}, {\em Birational equivalences and generalized weyl algebras}.
\newblock {\tt arXiv:2009.14801}, 2020.

\bibitem{marquis_introduction_2018}
{\sc T.~Marquis}, {\em An introduction to Kac-Moody groups over fields},
  Zürich, Switzerland, 2018.

\bibitem{McConnell:GKConjecture}
{\sc J.~C. McConnell}, {\em Representations of solvable {L}ie algebras and the
  {G}elfand-{K}irillov conjecture}, Proc. London Math. Soc. (3), 29 (1974),
  pp.~453--484.

\bibitem{MosinPanov:QuantumGK-I}
{\sc V.~G. Mosin and A.~N. Panov}, {\em Quotient skew fields and central
  elements of multiparametric quantizations}, Mat. Sb., 187 (1996), pp.~53--72.

\bibitem{Panov:quantumGelfandKirillov}
{\sc A.~Panov}, {\em Fields of fractions of quantum solvable algebras}, J.
  Algebra, 236 (2001), pp.~110--121.

\bibitem{Rosenber:NCAlgebraicGeometry}
{\sc A.~Rosenberg}, {\em Noncommutative algebraic geometry and representations
  of quantized algebras}, vol.~330 of Mathematics and its Applications, Kluwer
  Academic Publishers Group, Dordrecht, 1995.

\end{thebibliography}

\documentclass[12pt,a4paper]{amsart}

\usepackage[margin=25mm,top=28mm]{geometry}

\usepackage{amsmath, amssymb, amsgen, amsthm, amsfonts}
\usepackage{palatino}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc} 

\usepackage{hyperref}
\usepackage{color}

\newcommand{\G}[1]{\mathfrak{#1}}
\newcommand{\C}[1]{\mathcal{#1}}
\newcommand{\B}[1]{\mathbb{#1}}

\newcommand{\adlie}{\text{ad}}
\newcommand{\adgr}{\text{Ad}}

\renewcommand{\leq}{\leqslant}
\renewcommand{\geq}{\geqslant}

\numberwithin{equation}{section}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{corollary}[theorem]{Corollary}

\title{Birational Equivalences and Kac-Moody Algebras}

\author{Atabey Kaygun}
\address{Istanbul Technical University, Istanbul, Turkey.}
\email{kaygun@itu.edu.tr}

\begin{document}
\maketitle

\begin{abstract}
  We show that every Kac-Moody algebra is birationally equivalent to a smash biproduct of
  two copies of a Weyl algebra together with a polynomial algebra.  We also show that the
  same is true for quantized Kac-Moody algebras where one replaces Weyl algebras with their
  quantum analogues.
\end{abstract}


\section*{Introduction}

We show that every Kac-Moody algebra is birationally equivalent to a smash biproduct of two
copies of an algebra $\B{A}_{n-\ell,n}\otimes k[t_1,\ldots,t_\ell]$ defined as a product of
a Weyl algebra $\B{A}_{n-\ell,n}$ and a polynomial algebra $k[t_1,\ldots,t_\ell]$ where $n$
and $\ell$ are respectively determined by the size and the rank of the underlying
generalized Cartan matrix.  We also show that the same is true for the Drinfeld-Jimbo
quantization of a Kac-Moody algebra where one has to replace the Weyl algebra with its
quantum analogue.  The result we prove in this article is but a small step towards
reformulating the Gel'fand-Kirillov (GK) conjecture for Kac-Moody algebras and their
quantizations.

The GK-conjecture states that the universal enveloping algebra $U(\G{g})$ of a finite
dimensional Lie algebra $\G{g}$ is birationally equivalent to a Weyl algebra $\B{A}_n$ for
some $n$~\cite{GelfandKirillov:Conjecture-I, GelfandKirillov:Conjecture-II}.  The conjecture
is known to be false in general~\cite{AOV:CounterExamplesToGK}, but it is true for
$\G{gl}_n$, $\G{sl}_n$ and nilpotent Lie algebras~\cite{GelfandKirillov:Conjecture-I,
  GelfandKirillov:Conjecture-II}, for solvable Lie algebras~\cite{McConnell:GKConjecture,
  Joseph:GKConjecture}, and for every Lie algebra up to dimension
8~\cite{AOV:GKUptoDimension8}.  The corresponding conjecture for quantum deformations of Lie
algebras and Lie groups is known to be true in many cases, notably for $U_q(\G{sl}_2)$,
$\tilde{U}_q(\G{sl}_n)$, $\tilde{U}_q^{\geq 0}(\G{g})$, $\C{O}_q(M_n)$, and $\C{O}_q(G)$
when $q$ is not a root of unity, or is transcendental over
$\B{Q}$~\cite{AlevDumas:QuantumGK, FutornyHartwig:QuantumGK, MosinPanov:QuantumGK-I,
  Millet:QuantumGK, IoharaMalikov:QuantumGK, Caldero:quantumGelfandKirillov}, and all
$U_q^+(\G{g})$, and therefore, $U_q(\G{g})$ for solvable Lie algebras when $q$ is not a root
of unity~\cite{Panov:quantumGelfandKirillov}. A version of the conjecture has also been
studied by Colliot-Thélène, Kunyavski\u{ı}, Popov and Reichstein in~\cite{CKPR11} where the
authors investigated if algebra of functions on a Lie group $G$ or a Lie algebra $\G{g}$ is
purely transcendental over its invariant subalgebra.

The birational equivalence we prove in this article transforms the GK-conjecture for
Kac-Moody algebras and their quantizations to a similar question on smash biproducts of
(quantized) Weyl and polynomial algebras.  Inspired by the GK conjecture,
in~\cite{kaygun:GWAs} we conjectured that the universal enveloping algebra of a Lie algebra
$\G{g}$ is birationally equivalent to a smash product of a smooth algebra and a torus whose
rank is determined by the rank of the Cartan subalgebra of $\G{g}$. In this paper, we verify
that our conjecture is true for Borel algebras: any Borel subalgebra of a (quantized)
Kac-Moody algebra is birationally equivalent to the smash product of a polynomial algebra
and a torus.

\subsection*{Plan of the article}

Here is a plan of this article. In Section~\ref{sect:generalizedWeyl}, we define generalized
Weyl algebras bound by a generalized Cartan matrix.  We then show that when such an algebra
is of full rank (See Subsection~\ref{subsect:CartanDatum}) it is birationally equivalent to
the product of an ordinary Weyl algebra and a polynomial algebra. In
Theorem~\ref{thm:birational-UEA} we prove that the Borel subalgebras of Kac-Moody algebras
are birationally equivalent to a generalized Weyl algebra which in turn allows us to show
our main result in Corollary~\ref{cor:KacMoody}. Then we prove the analogous results for
Drinfeld-Jimbo quantizations of Kac-Moody algebras in Section~\ref{sect:quantizedKacMoody}.

\subsection*{Notation, conventions and some background}

\subsubsection*{Ground field}

Throughout the paper we fix a ground field $k$ of characteristic 0. All unadorned tensor
products are over $k$.

\subsubsection*{Ambient algebras}

All algebras are assumed to be unital and associative, but not necessarily commutative or
finite dimensional.  We also assume that our algebras are domains, i.e. devoid of any zero
divisors. There are several classes of algebras that we are going to use frequently in this
paper. These are as follows:
\begin{enumerate}

\item $k\{t_1,\ldots,t_n\}$ the noncommutative polynomial algebra on $n$ generators,

\item $k[t_1,\ldots,t_n]$ the commutative polynomial algebra on $n$ generators,

\item $\B{T}^n$ the Laurent polynomial algebra $k[t_1^{\pm 1},\ldots,t_n^{\pm 1}]$, and

\item $\B{A}_{m,n}$ the Weyl algebra given by the presentation
  \[ \frac{k\{x_1,\ldots,x_m,y_1,\ldots,y_n\}}{\langle[x_i,x_{i'}],\ [y_j,y_{j'}],\ [x_i,y_j] -
      \delta_{ij}\mid 1\leq i,i'\leq m,\ 1\leq j,j'\leq n\rangle}
  \]
  where $\delta_{uv}$ is the Kronecker delta.
\end{enumerate}

The classical Weyl algebra $\B{A}_n$ is $\B{A}_{n,n}$, and one has
\[ \B{A}_{m,n} \cong \B{A}_{\min(m,n)}\otimes
  k[t_1,\ldots,t_{|m-n|}] \] for every $m,n\in\B{N}$.

\subsubsection*{Ore sets and localizations}

A multiplicative submonoid $S$ of an algebra $A$ is called a \emph{right Ore set} if for
every $s\in S$ and $u\in A$ there are $s'\in S$ and $u'\in A$ such that $su=u's'$ . If $S$
is a right Ore set then one can invert the elements in $S$ to get an algebra $A_S$ and a
morphism of algebras $\iota_S\colon A\to A_S$ such that $\varphi(S)\subseteq A_S^\times$.

\subsubsection*{Birational equivalences}

In classical algebraic geometry, two irreducible algebraic variety $X$ and $Y$ are called
\emph{birationally equivalent} if there is a rational algebraic function $f\colon X\to Y$
that induces a bijection from a Zariski open subset $U$ of $X$ to a Zariski open subset $V$
of $Y$.  An affine (non-commutative) algebraic variety is called \emph{irreducible} if its
coordinate algebra has no idempotents other than 0 and 1.  With this notion at hand, we call
a morphism of (irreducible) affine non-commutative algebraic varieties
$\varphi\colon A\to A'$ as a \emph{birational equivalence} if there are suitable
localizations $A_S$ and $A'_{S'}$ such that that $\varphi(S)\subseteq S'$ and the extension
$\varphi_S\colon A_S\to A'_{S'}$ is an isomorphism of unital associative algebras.

\subsubsection*{Generalized Cartan matrices}

Throughout the article, we work with a fixed generalized symmetrizable Cartan matrix
$C = (a_{ij})$ of co-rank $\ell$, i.e. the null space of the matrix has dimension
$0\leq \ell\leq n$. 

\subsection*{Acknowledgements}

Most of this article was written while the author was on academic leave at Queen’s
University from Istanbul Technical University. The author would like to thank both
universities for their support.

\section{Generalized Weyl  Algebras}\label{sect:generalizedWeyl}

\subsection{Distributive laws}

Given two algebras $A$ and $B$, and a linear map $\rho\colon B\otimes A\to A\otimes B$ is
called a \emph{distributive law} if the vector space $A\otimes B$ is a unital associative
algebra with the multiplication $(\mu_A\otimes\mu_B)\circ (id_A\otimes\rho\otimes id_B)$ where
$\mu_A$ and $\mu_B$ are multiplication maps of the underlying algebras.  The product algebra is
called the \emph{smash biproduct} of $A$ and $B$, and is denoted by $A\#_\rho B$.  If the
distributive law is clear in the context, we may drop it from the notation.

\subsection{The ambient smash product}

Let us fix a commutative algebra $A$.  For a fixed $n\geq 1$, assume we have a collection of
pairwise commuting automorphisms $\sigma_i\in Aut(A)$ for every $1\leq i\leq n$.
Equivalently, we have a smash product algebra $A\# \B{T}^n$ given by the relations
$t_i a = \sigma_i(a)t_i$ for $a\in A$ and $1\leq i\leq n$.

\subsection{Generalized Weyl algebras}

Let us fix a sequence $\mathbf{b} = (b_1,\ldots,b_n)$ of elements from $A$.  Then we define
the \emph{generalized Weyl algebra of rank-$n$} $\B{W}_\sigma(A,\mathbf{b})$ as the
subalgebra of $A\# \B{T}^n$ generated by $A$ and elements of the form
\begin{equation}\label{eq:generators}
  b_i t_i^{-1}\quad\text{ and }\quad  -t_i
\end{equation}
for $1\leq i\leq n$.  See~\cite{Bavula:GWAI, Bavula:GlobalDim, Hodges:NCDeformations,
  Rosenber:NCAlgebraicGeometry}. See also~\cite{kaygun:GWAs} and references therein.  Now,
we extend \cite[Thm.2.1]{kaygun:GWAs} as follows:
\begin{proposition}\label{prop:birationalWeyl}
  $\B{W}_\sigma(A,\mathbf{b})$ is birationally equivalent to the smash product algebra
  $A\# \B{T}^n$.
\end{proposition}

\begin{proof}
  The subalgebra $\B{W}_\sigma(A,\mathbf{b})$ becomes the whole algebra $A\# \B{T}^n$ when
  we localize $\B{W}_\sigma(A,\mathbf{b})$ with respect to the Ore set generated by elements
  $\sigma_i^m(b_j^n)$ with $m\in\B{Z}$, $n\in\B{N}$ and $i,j=1,\ldots,n$.
\end{proof}

\section{Cartan Datum}\label{sect:CartanDatum}

In this Section we assume $\B{W}_\sigma(A,\mathbf{b})$ is a generalized Weyl algebra.

\subsection{Twisted differentials}

Let us use $D_i$ for the discrete total derivative of the automorphism $\sigma_i$ for every
$1\leq i\leq n$.  These are linear operators acting on $A$ as
\[ D_i(f) = t_i f t_i^{-1} - f = \sigma_i(f) - f \] for every $f\in A$ and $1\leq i\leq n$.
Observe that $D_i$ is as a right, or equivalently a left, $\sigma_i$-derivation since
\[ D_i(fg) = \sigma_i(fg)-fg = 
  D_i(f)\sigma_i(g) + f D_i(g) \] for every $f,g\in A$ and $1\leq i\leq n$.

\subsection{Cartan datum}\label{subsect:CartanDatum}

Let us fix a generalized Cartan matrix $C=(a_{ij})$.  A generalized Weyl algebra
$\B{W}_\sigma(A,\mathbf{b})$ is said \emph{to be bound by $C$} if the collection
$\mathbf{b}=(b_1,\ldots,b_n)$ satisfies the following conditions
\begin{align}
  D_iD_j(b_j)  & = a_{ji} \quad\text{ for every $i$ and $j$}, \label{eq:CS2}\\
  D_i^{1-a_{ij}}(b_j) & = 0 \quad\text{ for every } i\neq j.\label{eq:CS1}
\end{align}
We will say that a generalized Weyl algebra bound by $C$ \emph{has full rank} when the
elements $D_i(b_i)$ are algebraically independent and the subalgebra they generate is
birationally equivalent to $A$.  We are going to use $\B{W}_\sigma(A,C)$ to denote a
generalized Weyl algebra $\B{W}_\sigma(A,\mathbf{b})$ bound by $C$, and we are going to
refer to it as \emph{the Cartan datum} when the algebra has full rank.

\subsection{Cartan datum and ordinary Weyl algebras}

Given a linear endomorphism $T\colon V\to V$ of a finite dimensional vector space $V$, one
can split $V$ into direct sum of two spaces $V_1\oplus V_2$ where $T$ is identically zero on
$V_2$ and invertible on $V_1$.  Thus one can write a new invertible endomorphism
$Q\colon V\to V$ such that $QT = id_{V_1}\oplus 0_{V_2}$.  We call the matrix $Q$ as the
\emph{quasi-inverse} of $T$.

\begin{proposition}\label{prop:embeddingWeyl}
  Assume $C$ has co-rank $\ell$, and let $\B{W}_\sigma(A,C)$ has full rank.  Then
  $\B{W}_\sigma(A,C)$ is birationally equivalent to the product algebra
  $\B{A}_{n-\ell,n}\otimes k[x_{n-\ell+1},\dots,x_n]$.
\end{proposition}

\begin{proof}
  Assume $C=(a_{ij})$ has rank $n-\ell$ and let $Q = (c_{ij})$ be its quasi-inverse.  Then
  we have
  \[ \sum_{u=1}^n c_{ju}a_{ui} =
    \begin{cases}
      1 & \text{ if } 1\leq i\leq n-\ell \text{ and } i=j\\
      0 & \text{ otherwise.}
    \end{cases}
  \] Now, we let $h_i = D_i(b_i)$ and define $\alpha_i$ in $\B{W}_\sigma(A,C)$ given by
  \begin{equation}
    \label{eq:alternate-basis}
    \alpha_i = \sum_{u=1}^n c_{iu} h_u.
  \end{equation}
  These elements satisfy
  \begin{equation}
    \label{eq:differentials}
    D_i (\alpha_j) = \sum_u c_{ju}D_i(h_u) = \sum_u c_{ju}a_{ui} = 
    \begin{cases}
      1 & \text{ if } 1\leq i\leq n-\ell \text{ and } i=j\\
      0 & \text{ otherwise }
    \end{cases}
  \end{equation}
  This means $\alpha_i$ is in the centre when $n-\ell+1\leq i\leq n$. Then we have a
  morphism of algebras $\B{A}_{n-\ell,n}\otimes k[x_{n-\ell+1},\ldots,x_n]\to A\#\B{T}^n$
  given by
  \[ x_i\mapsto \alpha_i t_i^{-1} \qquad y_i \mapsto -t_i. \] One can
  check that it is well-defined since defining relations are
  satisfied.  Since the $Q$ is invertible, the image subalgebra is
  isomorphic to the subalgebra generated by $h_it_i^{-1}$ and
  $t_i$. Moreover, since the algebra has full rank, the elements $h_i$
  birationally generate $A$ as a polynomial algebra. If we localize
  $A\# \B{T}^n$ on the Ore set generated by elements $h_i$ we get the
  desired birational equivalence between
  $\B{A}_{n-\ell,n}\otimes k[x_{n-\ell+1},\ldots,x_n]$ and
  $A\#\B{T}^n$. Now, we use Proposition~\ref{prop:birationalWeyl}.
  The result follows.
\end{proof}

\section{Birational Equivalence for Kac-Moody Algebras}

\subsection{Kac-Moody algebra of a generalized Cartan matrix}

After~\cite[Defn. 3.17]{marquis_introduction_2018}, we define the Kac-Moody Lie algebra
$\G{g}(C)$ associated with the generalized Cartan matrix $C=(a_{ij})$ as the Lie algebra
generated by vectors $E_i, F_i, H_i$ for $1\leq i\leq n$ subject to the relations
\begin{align}
  [H_i,H_j] = & 0, &          [E_i, F_j] = & \delta_{ij} H_j,   \label{eq:UEA1}\\
  [H_i,E_j] = & a_{ij} E_j, &   [H_i, F_j] = & -a_{ij} F_j, \label{eq:UEA2}
\end{align}
for every $1\leq i,j\leq n$, and
\begin{align}
  \adlie(E_i)^{1-a_{ij}}(E_j) = & 0 &   \adlie(F_i)^{1-a_{ij}}(F_j) = & 0 \label{eq:UEA3}
\end{align}
for every $i\neq j$ where $\adlie(x)(y)$ is defined to be $xy-yx$ for every
$x,y\in \G{g}(C)$.

The Kac-Moody algebra of a generalized Cartan matrix $C$ is the universal enveloping algebra
of the Kac-Moody Lie algebra $\G{g}(C)$.  We will use $U(\G{g})$ for the algebra
$U(\G{g}(C))$ dropping $C$ from the notation.

\subsection{Merging subalgebras}\label{sect:smashProduct}

The following subalgebras of $U(\G{g})$ are going to be used in the sequel:
\begin{enumerate}

\item The subalgebra $U^0(\G{g})$ generated by $H_i$ for $1\leq i\leq n$.

\item The subalgebra $U^{\geq 0}(\G{g})$ generated by $E_i$ and $H_i$ for $1\leq i\leq n$.

\item The subalgebra $U^{\leq 0}(\G{g})$ generated by $F_i$ and $H_i$ for $1\leq i\leq n$.
  
\end{enumerate}

Consider the algebras $U^{\geq 0}(\G{g})$ and $U^{\leq 0}(\G{g})$, and define a distributive
law of the form
\[ \omega\colon U^{\leq 0}(\G{g})\otimes_{U^0(\G{g})} U^{\geq 0}(\G{g}) \to U^{\geq
    0}(\G{g})\otimes_{U^0(\G{g})} U^{\leq 0}(\G{g}) \] given by
\[ \omega(F_j E_i) = E_i F_j - \delta_{ij} H_i \] for every $1\leq i,j\leq n$.  Since we have a
Poincare-Birkhoff-Witt basis, one can easily see that the resulting product is exactly
$U(\G{g})$.

\subsection{The canonical Cartan datum of a Kac-Moody algebra}

Let $A$ be the polynomial algebra $k[h_1,\ldots,h_n]$ and let us define
\begin{equation}
  \label{eq:2}
  \sigma_i(h_j) = h_j + a_{ji} 
\end{equation}
for every $1\leq i,j\leq n$.  Notice that with this choice we get $D_i(h_j) = a_{ji}$ for
every $i$ and $j$. Let us recall the elements $\alpha_i$ from~\eqref{eq:alternate-basis} and
we choose
\begin{equation}
  \label{eq:4}
  b_i = \frac{1}{4}h_i(h_i - 2) + \beta_i
\end{equation}
for every $1\leq i\leq n$ where $\beta_i\in k[h_1,\ldots,h_n]$ is an undetermined element in
the subalgebra generated by $\alpha_j$ with $j\neq i$ that we constructed in the proof of
Proposition~\ref{prop:embeddingWeyl}.  We immediately get that $D_i(b_i) = h_i$ for every
$1\leq i\leq n$. Moreover, we see that
\[ D_iD_j(b_j) = D_i(h_j) = a_{ji}. \] Since in the basis $\alpha_i$'s the twisted
differentials $D_i$'s behave like ordinary differential operators, one can solve the system
of differential equations $D^{1-a_{ij}}(b_j)$ for $b_i$'s and determine $\beta_i$'s. This
means the generalized Weyl algebra $\B{W}_\sigma(A,\mathbf{b})$ is bound by $C$ and has
full-rank, i.e. we have a Cartan datum $\B{W}_\sigma(A,C)$.


\subsection{Birational equivalence for Kac-Moody algebras}

\begin{theorem}\label{thm:birational-UEA}
  The algebras $U^{\geq 0}(\G{g})$ and $U^{\leq 0}(\G{g})$
  are birationally equivalent to the product $\B{A}_{n-\ell,n}\otimes k[t_1,\ldots,t_\ell]$
  when the underlying Cartan matrix has co-rank $\ell$.
\end{theorem}

\begin{proof}
  We will give the proof for $U^{\geq 0}(\G{g})$. The proof for $U^{\leq 0}(\G{g})$ is
  similar.  

  We have a well-defined morphism of algebras of the form
  $U^{\geq 0}(\G{g})\to \B{S}_\sigma(k[h_1,\ldots,h_n],C)$ given by
  \begin{equation}
    \label{eq:13}
    H_i \mapsto h_i \qquad E_i \mapsto \left(\frac{1}{4}h_i(h_i-2) + \beta_i\right) t_i^{-1}
  \end{equation}
  for $1\leq i\leq n$ which is a birational equivalence when we invert $H_i$ and $h_i$'s and
  $E_i$ and $e_i$'s.  On the other hand Proposition~\ref{prop:embeddingWeyl} tells us that
  $\B{W}_\sigma(k[h_1,\ldots,h_n] ,C)$ is birationally equivalent to
  $\B{A}_{n-\ell,n}\otimes k[x_{n-\ell+1},\ldots,x_n]$ since
  $\B{W}_\sigma(k[h_1,\ldots,h_n] ,C)$ has full rank.  The result follows.
\end{proof}

\begin{corollary}\label{cor:KacMoody}
  Let $C$ be a generalized Cartan matrix of co-rank $\ell$ and let $\B{W}_\sigma(A,C)$ be a
  Cartan datum.  The Kac-Moody algebra associated with $C$ is birationally equivalent to the
  smash biproduct of two copies of the algebra $\B{W}_\sigma(A,C)$, or equivalently smash
  biproduct of two copies of the algebra $\B{A}_{n-\ell,n}\otimes k[t_1,\ldots,t_\ell]$.
\end{corollary}

\begin{proof}
  Follows Proposition~\ref{prop:embeddingWeyl} and Theorem~\ref{thm:birational-UEA}.
\end{proof}

\section{Birational Equivalence for Quantized Kac-Moody algebras}\label{sect:quantizedKacMoody}

As before, let us fix an algebra $A$, a set of commuting automorphisms $\sigma_i\in Aut(A)$ and
elements $b_i\in A$ for $1\leq i\leq n$.  This time we assume that the Cartan matrix is
symmetrizable, i.e. there are scalars $d_i\in k^\times$ such that $d_i a_{ij} = d_j a_{ij}$ for
$1\leq i,j\leq n$.

\subsection{Quantized Cartan datum}

A generalized Weyl algebra $\B{W}_\sigma(A,\mathbf{b})$ is called the \emph{quantum
  generalized Weyl algebra} bound by a generalized Cartan matrix $C = (a_{ij})$ if the
elements $b_1,\ldots,b_n$ satisfy the following conditions
\begin{align}
  \sigma_i(b_j) = & q^{d_i a_{ij}} b_j\ \text{ for every $i$ and $j$,}   \label{eq:qCS2}\\
  D_i^{[1-a_{ij}]}(b_j) = & 0\ \text{ for every } i\neq j \label{eq:qCS1}
\end{align}
where we define
\begin{equation}
  \label{eq:6}
  D_i^{[m]} = \prod_{\ell=0}^{m-1}(\sigma_i - q^{2\ell d_i})
\end{equation}
for every $m\geq 0$ and $1\leq i\leq n$.  As before, we say $\B{W}_\sigma^q(A,C)$ is of full
rank if the subalgebra generated by $b_i$'s are algebraically independent and birationally
generate $A$. We are going to refer such an algebra as \emph{the quantum Cartan datum}, and
denote it by $\B{W}^q_\sigma(A,C)$.

\subsection{Quantized Cartan datum and quantized Weyl  algebras}\label{sect:quantized-Weyl-vs-Borel}

Let $Q = (c_{ij})$ be the quasi-inverse of the Cartan matrix, and let $g_i$ be the smallest
positive integer such that $c_{ij}g_j$ is an integer for every $1\leq i,j\leq n$.  We define
the quantum Weyl algebra $\B{A}^q_{m,n}$ with the presentation
\begin{equation}
  \label{eq:11}
  \frac{k\{x_1,\ldots,x_m,y_1,\ldots,y_n\}}{\langle [x_i, x_{i'}], [y_j, y_{j'}], y_j x_i - q^{g_i\delta_{ij}}x_i y_j\mid  i,i'=1,\ldots,m,\ j,j'=1,\ldots,n\rangle}.
\end{equation}

\begin{proposition}\label{prop:rational-Weyl}
  Assume $C$ has co-rank $\ell$. Then the quantized Cartan datum $\B{W}_\sigma^q(A,C)$ is
  birationally equivalent to $\B{A}^q_{n-\ell,n}\otimes k[t_1,\ldots,t_\ell]$.
\end{proposition}

\begin{proof}
  Let $S$ be the Ore set in $A\# \B{T}^n$ generated by the elements $b_i$ for
  $1\leq i\leq n$.  For each $1\leq i\leq n$ we consider
  \begin{equation}
    \label{eq:10}
    \omega_i := \prod_{u=1}^n b_u^{c_{ui}g_i}.
  \end{equation}
  in $(A\# \B{T}^n)_S$.  Now, observe that
  \[ \sigma_j(\omega_i) = \prod_u \sigma_j (b_u^{c_{ui}g_i}) = \prod_u
    q^{a_{ju}c_{ui}g_i}b_u^{c_{ui}g_i} =
    \begin{cases}
      q^{g_i}\omega_i & \text{ if } 1\leq i=j\leq n-\ell\\
      \omega_i & \text{ otherwise.}
    \end{cases}
  \] The elements $\omega_i$ need not be elements in $A$ since we may have negative exponents.
  This problem is solved by using our suitable localization.  Now, one can define an algebra
  map after such a localization $\B{A}_{n-\ell,n}^q\to (A\# \B{T}^n)_S$ by
  \begin{equation}
    \label{eq:quantum-weyl-map}
    x_i \mapsto \omega_it_i^{-1} \qquad y_i\mapsto t_i 
  \end{equation}
  for $1\leq i\leq n$.  One can check that
  the defining relations are satisfied. Now, determining whether this embedding is a
  birational equivalence boils down to checking if the embedding of the subalgebra of $A$
  generated by $\omega_1,\ldots,\omega_n$ is a birational equivalence. The result follows.
\end{proof}

\subsection{Quantized Kac-Moody algebra}\label{sect:quantumKacMoody}

We fix a $q\in k^{\times}$ and let $d_1,\ldots,d_n$ be the set of positive integers such
that $d_i a_{ij} = d_j a_{ji}$ for every $1\leq i,j\leq n$.  We define the quantum Kac-Moody
algebra $U_q(\G{g})$ associated with the generalized Cartan matrix as the algebra generated
by non-commuting indeterminates $E_i, F_i, K_i^{\pm 1}$ for $1\leq i\leq n$ subject to the
following relations
\begin{align}
  K_iK_j = & K_jK_i, &
 [E_i, F_j] = & \delta_{ij} \frac{K_i - K_i^{-1}}{q^{d_i}-q^{-d_i}},   \label{eq:qUEA1}\\
  E_j K_i = & q^{-d_i a_{ij}} K_i E_j, &   F_j K_i = & q^{d_i a_{ij}} K_i F_j, \label{eq:qUEA2}
\end{align}
for every $1\leq i,j\leq n$ and
\begin{align}
  \adlie_q(E_i)^{1-a_{ij}}(E_j) = & 0 &   \adlie_q(F_i)^{1-a_{ij}}(F_j) = & 0 \label{eq:qUEA3}
\end{align}
where $\adlie_q(E_i)(E_j)$ and $\adlie_q(F_i)(F_j)$ are defined to be
\begin{equation}
  \label{eq:17}
  \adlie_q(E_i)(E_j) = E_i E_j - q^{d_i a_{ij}} E_j E_i\quad\text{ and }\quad
  \adlie_q(F_i)(F_j) = F_i F_j - q^{-d_i a_{ij}} F_j F_i  
\end{equation}
for every $i\neq j$.  

\subsection{Localization of a quantized Kac-Moody algebra}

If we localize $U_q(\G{g})$ with respect to $E_i$ and $F_i$ for $1\leq i\leq n$, the
defining relation~\eqref{eq:qUEA2} manifests itself as
\begin{equation}
  \label{eq:localized-qCS2}
  \adgr(K_j^{-1}E_j)(K_i) = q^{-d_i a_{ij}}K_i\qquad \adgr(K_j^{-1}F_j)(K_i) = q^{d_i a_{ij}}K_i
\end{equation}
and the Chevalley-Serre relations become
\begin{equation}
  \label{eq:localized-qCS1}
  \prod_{\ell=0}^{-a_{ij}}(\adgr(K_i^{-1}E_i)-q^{2\ell d_i})(E_j) = 0\qquad
  \prod_{\ell=0}^{-a_{ij}}(\adgr(K_i^{-1}F_i)-q^{2\ell d_i})(F_j) = 0
\end{equation}
for every $1\leq i,j\leq n$.

\subsection{Merging quantized Borel subalgebras}\label{sect:quantum-merging}

As in the case of Kac-Moody algebras, we have following subalgebras.
\begin{enumerate}
\item The Cartan subalgebra $U_q^0(\G{g})$ generated by $K_i^{\pm 1}$ for $1\leq i\leq n$.
\item The Borel subalgebra $U_q^{\geq 0}(\G{g})$ generated by $E_i$ and $K_i^{\pm 1}$ for
  $1\leq i\leq n$.
\item The Borel subalgebra $U_q^{\leq 0}(\G{g})$ generated by $F_i$ and $K_i^{\pm 1}$ for
  $1\leq i\leq n$.
\end{enumerate}

Again as before, one can merge the Borel subalgebras $U_q^{\geq 0}(\G{g})$ and
$U_q^{\leq 0}(\G{g})$ via a suitable distributive law:
\[ \omega_q\colon U_q^{\leq 0}(\G{g})\otimes_{U_q^0(\G{g})} U_q^{\geq 0}(\G{g}) \to
  U_q^{\geq 0}(\G{g})\otimes_{U_q^0(\G{g})} U_q^{\leq 0}(\G{g}) \] given by
\[ \omega(F_i E_j) = E_j F_i - \delta_{uv}\frac{K_i-K_i^{-1}}{q^{d_i}-q^{-d_i}} \] for every
$1\leq i,j\leq n$.  Since we have a Poincare-Birkhoff-Witt basis, one can easily see that
the resulting product is exactly $U_q(\G{g})$.

\subsection{Birational equivalences for quantum Kac-Moody algebras}\label{sect:birational-qCS}

\begin{theorem}\label{thm:quantumBorel}
  The quantum Borel algebras $U_q^{\geq 0}(\G{g})$ and $U_q^{\leq 0}(\G{g})$ are birationally
  equivalent to the product algebra $\B{A}^q_{n-\ell,n}\otimes k[t_1,\ldots,t_\ell]$ when the
  underlying Cartan matrix has co-rank~$\ell$.
\end{theorem}

\begin{proof}
  Let $A$ be the torus $\B{T}^n$ generated by $K_1,\ldots,K_n$ and let
  \begin{equation}
    \label{eq:18}
    \sigma_i(K_j) = q^{-d_i a_{ij}} K_j
  \end{equation}
  for every $1\leq i,j\leq n$.  If we let $b_i=K_i^{-1}$ then the defining
  relation~\eqref{eq:qCS2} is satisfied in its localized form~\eqref{eq:localized-qCS2} for
  every $1\leq i,j\leq n$.  The defining equation~\eqref{eq:qCS1} is satisfied because we
  have~\eqref{eq:localized-qCS1}.  This means the algebra isomorphism
  $U_q^{\geq 0}(\G{g})\to \B{W}_\sigma^q(\B{T}^n,C)$ given by
  \[ K_i \mapsto K_i,\qquad E_i \mapsto K_i^{-1} t_i^{-1} \] for every $1\leq i\leq n$ is
  well-defined after a suitable localization.  One can also see that
  $\B{W}_\sigma^q(\B{T}^n,C)$ is full-rank since we use $b_i = K_i$ for $1\leq i\leq n$.  So,
  we obtain a birational equivalence of the form
  $U^{\geq 0}_q(\G{g})\to \B{W}_\sigma^q(\B{T}^n,C)$.  We also have another birational
  equivalence of the form
  $\B{A}^q_{n-\ell,n}\otimes k[x_{n-\ell+1},\ldots,x_n]\to \B{W}_\sigma^q(\B{T}^n,C)$ by
  Proposition~\ref{prop:rational-Weyl}.  The result follows.
\end{proof}

\begin{corollary}\label{cor:quantumKacMoody}
  The quantized Kac-Moody algebra $U_q(\G{g})$ is birationally equivalent to a smash biproduct
  of two copies of the product algebra $\B{A}^q_{n-\ell,n}\otimes k[t_1,\ldots,t_\ell]$ when
  the underlying Cartan matrix has co-rank $\ell$.
\end{corollary}

\begin{proof}
  Follows from Proposition~\ref{prop:rational-Weyl}, Theorem~\ref{thm:quantumBorel}, and
  Section~\ref{sect:quantum-merging}.
\end{proof}

\bibliographystyle{siam}
\bibliography{references}{}

\end{document}